### v2

@@ 3.18 Masking Strategy, Model Architecture, Downstream Task and Objective Function
[103] developed a lightweight MAE tailored for video anomaly detection. This MAE model leveraged weights derived from motion gradients to emphasize foreground objects in the reconstruction loss. Additionally, they enhanced the training procedure by introducing synthetic anomalies, while using normal frames as the target for the reconstruction loss. Later stages of training involve a student decoder, which learns to mimic the output of the main (teacher) decoder, further refining the detection process.
@@ 6 Performance Overview
Other significant observations, based on the results presented in Table 4, refer to the masking strategy and the types of features that are reconstructed. First, in videos, the temporal correlation between frames [35] can introduce a bias in the model, allowing it to reconstruct masked patches by leveraging information from adjacent frames. To alleviate this issue, OmniMAE [38] randomly masks patches from the videos similar to MAE [9], but uses a higher masking ratio, $95\%$ instead of $75\%$. However, this masking strategy performs poorly compared with those used in VideoMAE [35] and VideoMAEv2 [57], which are designed to mask temporal tubes instead of spatial patches. The latter approach is more effective in preventing information leakage from adjacent frames. Second, the top performing methods, such as MVD [183] and EVA [75], benefit from the use of high-level features as target features for reconstruction. These features help address the tendency of pixel-level MIM models to focus on spatial and structural information, leading to improved performance on downstream tasks.
@@ 7 Closing Remarks and Future Directions
Another promising direction for masked image modeling is represented by the masking strategy. While a random strategy may perform well, it has been demonstrated that an informed masking policy that focuses on hiding the salient information of the input is superior. Thus, future endeavors may attempt to formulate various guided masking strategies, some of which can be specific to the downstream task.
Some papers approached masked image modeling either through a reconstruction or a contrastive objective, and even both. These two pre-training strategies could be further extended, or even combined with other methods for better performance. Furthermore, another promising research direction consists of studying the target features on which the pre-training is applied.
While many papers focused on images as the only input modality, some promising results were shown on multimodal inputs. Each modality holds a different type of information, and these can be combined to learn richer feature representations. Distilling the knowledge of each modality into a single network could represent a stepping stone in artificial intelligence.
@@
Funding. This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906.

### v3

@@ 3.18 Masking Strategy, Architecture, Task and Objective
[120] developed a lightweight MAE tailored for video anomaly detection. This MAE model leveraged weights derived from motion gradients to emphasize foreground objects in the reconstruction loss. Additionally, the authors enhanced the training procedure by introducing synthetic anomalies, while using normal frames as the target for the reconstruction loss. Later stages of training involve a student decoder, which learns to mimic the output of the main (teacher) decoder, further refining the detection process.
@@ 6 Performance Overview
Other significant observations, based on the results presented in Table 4, refer to the masking strategy and the types of features that are reconstructed. First, in videos, the temporal correlation between frames [51] can introduce a bias in the model, allowing it to reconstruct masked patches by leveraging information from adjacent frames. To alleviate this issue, OmniMAE [54] randomly masks patches from videos, similar to MAE [9], but uses a higher masking ratio, $95\%$ instead of $75\%$. However, this masking strategy performs poorly compared with those used in VideoMAE [51] and VideoMAEv2 [73], which are designed to mask temporal tubes instead of spatial patches. The latter approach is more effective in preventing information leakage from adjacent frames. Second, the top performing methods, such as MVD [201] and EVA [93], benefit from the use of high-level features as target features for reconstruction. These features help address the tendency of pixel-level MIM models to focus on spatial and structural information, leading to improved performance on downstream tasks.
@@ 7 Closing Remarks and Future Directions
While masked image modeling (MIM) has achieved remarkable progress in self-supervised visual representation learning, several promising directions remain open for future exploration. One important area is the extension of MIM to multi-modal and cross-modal learning scenarios, such as integrating vision with language or audio signals, which could further enhance contextual understanding. Each modality holds a different type of information, and these can be combined to learn richer feature representations. Distilling the knowledge of each modality into a single network could represent a stepping stone in artificial intelligence.
Another avenue is the development of more efficient masking strategies and reconstruction objectives that better align with downstream tasks beyond classification, such as segmentation, detection, or 3D understanding. While a random strategy may perform well, it has been demonstrated that an informed masking policy that focuses on hiding the salient information of the input is superior [67]. Thus, future endeavors may attempt to formulate various guided masking strategies, some of which can be specific to the downstream task. Additionally, there is growing interest in making MIM more resource-efficient.
Another future direction is the integration of physical priors or domain-specific laws into the MIM framework. For domains where data is governed by known physical constraints (e.g. medical imaging, climate science, robotics), embedding known physical constraints or conservation laws into the masking or reconstruction process could guide the model toward more meaningful and generalizable representations.
Finally, understanding the theoretical underpinnings of why and how MIM works so effectively remains an open challenge. Insights on this path could drive the design of next-generation self-supervised learning algorithms.
@@
Funding. This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906. This work was also supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P7-7.1-PED-2024-1856, within PNCDI IV.